The Market. What is the competitive market equilibrium?
D: P = 220 - 4Q
S: P = 40 + 2Q
1
Expert's answer
2016-05-19T12:40:02-0400
The market equilibrium occurs when the quantity of good supplied is equal to quantity of goods demanded in a market.
D=S
220-4Q=40+2Q
Q =30

If equilibrium quantity equals 30, we can plug the equilibrium quantity into ether demand or supply function and receive market equilibrium price.
P=220-4*30=100
P= 40 + 2*30=100

The equilibrium quantity equals 30, the equilibrium price equals 100.